# Quantum Semantic Communications for Resource-Efficient Quantum Networking

### v2

@@ II-A Quantum Data Embedding
where $|{a_{0}}|^{2}+|{a_{1}}|^{2}+...+|{a_{d_{1}-1}}|^{2}=1$. The quantum state |$\ket{\varphi}$ can also be represented within a density matrix as a sum of outer products of basis states, i.e., $\rho\equiv\sum_{i=0}^{d_{1}-1}p_{i}\ket{i}\bra{i}$ |, where $p_{i}$ represents the probability of the state being in the |$\ket{i}$ state. The von Neumann entropy can be defined as [12], $\mathbb{H}_{q}=\mbox{Tr}\left(\rho\ln\rho\right)$, where “Tr" denotes the trace of a matrix. After pre-processing our classical data by applying contrastive learning to disentangle semantically-rich data points from semantically-poor data points [15], we then encode the classical data ${\boldsymbol{x}}\in\mathcal{X}$ into quantum states |$\ket{\psi({\boldsymbol{x}})}=\ket{y}$ using a quantum feature map $\psi:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{Y}\subset\mathcal{H}_{d_{1}}$. Here, the Hilbert space’s dimension, $d_{1}$, is much larger than $N$, the dimension of the data in $\mathcal{X}$. This is due to the potential lack of linear separability in the raw data space $\mathcal{X}$ for the data points ${\boldsymbol{x}}$. Consequently, transforming it into a higher-dimensional space $\mathcal{Y}$ becomes essential, facilitating the clustering of data points based on the semantic concepts, as described later. The quantum feature map $\psi$, which maps ${\boldsymbol{x}}\rightarrow\ket{y}$, is implemented via a quantum circuit $U_{\psi,d_{1}}({\boldsymbol{x}})$, called the quantum-embedding circuit. This circuit feeds the classical data ${\boldsymbol{x}}$ to quantum gates applied on ground states, |$\ket{0_{d_{1}}}$, in $\mathcal{H}_{d_{1}}$. To ensure computational feasibility and to avoid exponentially growing Hilbert spaces, quantum embedding circuits in the QSC framework are designed such that they compromise between the number of encoding qudits, and the dimensionality, $d_{1}$, of the Hilbert space. This optimization is mainly dependent on the size and dimensionality of the classical data, along with the available quantum computing powers. Essentially, such a quantum feature map yields the quantum embedded states |$\ket{y}$ according to: $U_{\psi,d_{1}}({\boldsymbol{x}})\ket{0_{d_{1}}}=\ket{y}$ [16]. Subsequently, these quantum states are stored in tailored structures. After embedding the data into quantum states, the next step is apply QML to construct efficient quantum semantic representations, as described next.
@@ II-B1 Problem formulation for quantum k-means clustering
where $\mathbb{H}(c\mid\ket{y})=\sum\limits_{i}\sum\limits_{k}p(c_{k}\mid\ket{y_{i}})\ln p(c_{k}\mid\ket{y_{i}})$. The penalty term $\mathbb{I}(\ket{y};c)$ ensures that the semantic information conveyed by the $k$-means clustering that identifies the semantic concepts is maximized.
@@ II-B2 Solution to (2)
Here, the stored quantum vectors are initialized to different clusters either in an arbitrary fashion or by utilizing efficient heuristic approaches. Then, multiple iterations are performed such that in each iteration, the goal is to minimize the loss function in (2), which ensures that each vector is assigned to the cluster with the closest centroid. This process ensures that semantically similar data vectors are clustered together (as guaranteed by the penalty term in (2)). This clustering aids in uncovering hidden patterns that might remain undiscovered in the raw classical data without the utilization of . All the quantum states belonging to the same cluster form a semantic space and are semantically similar, as each centroid represents a semantic concept. Henceforth, the set $\Phi$ of |$\ket{\phi_{i}},\forall i$ forms a common language, composed of the syntactic space [13]. Along the lines of [13], the amount of semantic information conveyed by any cluster $c_{k}$ can be written as the average across of the information represented by all the elements in the cluster $c_{i}$:
$\mathbb{S}(c_{k})=\sum\limits_{\ket{y}\in\mathcal{Y}}p(\ket{y})\log\frac{p(\ket{y}\mid c_{k})}{p(\ket{y})}.$ (3)
Given that the number of semantically distinct concepts $K\leq\lvert\mathcal{X}\rvert$, we chose to encode the semantic concepts into semantic representations |$\ket{s}\in\mathcal{S}$ that exists in a lower dimensional Hilbert space for quantum communication $\mathcal{H}_{d_{2}}$, where $d_{2}\leq d_{1}$. Semantic representation ensures that all the concepts that are identical from the receiver’s perspective are encoded with the same representation that gets communicated over the QCN, thus saving significant amount of quantum communication resources. A semantic representation is defined as the mapping, $f:\mathcal{Y}\,\cup\,\Phi\rightarrow\mathcal{S}$ and is implemented as a probabilistic mapping $p(\ket{s_{i}}\mid c_{k})$. The number of quantum communication resources $C=\lvert\mathcal{S}\rvert$ corresponds to the number of quantum states transmitted, which can be less than the number of clusters $K$ as the semantic representation mapping $f$ consolidates all semantically similar concepts and represents them using the same |$\ket{s_{i}}\in\mathcal{S}$.
@@ III QSC Problem Formulation and Performance Assessment Framework
First, regarding the quantum communication aspect, today’s quantum computers are devices that incorporate noise and various unavoidable losses that affect the quantum circuits employed in quantum embedding, entanglement generation, and entanglement measurement operations. A worst-case general quantum channel model that captures such noise sources in QCNs is the quantum depolarizing noise model, where a quantum state $\rho$ is preserved with a probability $1-\lambda$ and lost with probability $\lambda$, where $\lambda$ is a real-valued parameter [19]. This model, denoted as $\Delta_{\lambda}$, maps a quantum state $\rho\in\mathcal{H}_{d}$ into a linear combination that comprises $\rho$ itself, and the $d\times d$ identity matrix $\boldsymbol{I}_{d}$. It is a trace-preserving, completely positive map, given by:
$\Delta_{\lambda}(\rho)=(1-\lambda)\rho+\frac{\lambda}{d}\boldsymbol{I}_{d}.$ (7)
It is important to note that parameter $\lambda$ is bounded by: $-\frac{1}{d^{2}-1}\leq\lambda\leq 1,$ so as to guarantee the complete positivity condition. The quality of the communicated quantum resources is quantified using the quantum communication fidelity, $F_{c}$, of the considered received quantum states which captures the unavoidable noise that every quantum state undergoes when transmitted over quantum channels. In this regard, when the generated entangled pairs of qudits are considered to have maximally entangled measurements and channels, the average quantum communication fidelity that corresponds for the depolarizing noise will be [19]:
$\vspace{-1mm}\small\langle F_{c}\rangle=1-\frac{d-1}{d}\lambda,$ (8)
where $d$ is the dimension of the qudits. If a low fidelity is encountered, entanglement purification protocols may be applied [20] to achieve a desired minimum quality of the entanglement connection. Meanwhile, when the entanglement connection between the transmitter and receiver is characterized with a high fidelity, then the semantic-representing quantum states will be accurately received.

### v3

@@ II-A Quantum Data Embedding
where $|{a_{0}}|^{2}+|{a_{1}}|^{2}+...+|{a_{d_{1}-1}}|^{2}=1$. The quantum state |$\ket{\varphi}$ can also be represented within a density matrix as a sum of outer products of basis states, i.e., $\rho\equiv\sum_{i=0}^{d_{1}-1}p_{i}\ket{i}\bra{i}$ |, where $p_{i}$ represents the probability of the state being in the |$\ket{i}$ state. The von Neumann entropy can be defined as [12], $\mathbb{H}_{q}=-\mbox{Tr}\left(\rho\ln\rho\right)$, where “Tr" denotes the trace of a matrix. After pre-processing our classical data by applying contrastive learning to disentangle semantically-rich data points from semantically-poor data points [15], we then encode the classical data ${\boldsymbol{x}}\in\mathcal{X}$ into quantum states |$\ket{\psi({\boldsymbol{x}})}=\ket{y}$ using a quantum feature map $\psi:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{Y}\subset\mathcal{H}_{d_{1}}$. Here, the Hilbert space’s dimension, $d_{1}$, is much larger than $N$, the dimension of the data in $\mathcal{X}$. This is due to the potential lack of linear separability in the raw data space $\mathcal{X}$ for the data points ${\boldsymbol{x}}$. Consequently, transforming it into a higher-dimensional space $\mathcal{Y}$ becomes essential, facilitating the clustering of data points based on the semantic concepts, as described later. The quantum feature map $\psi$, which maps ${\boldsymbol{x}}\rightarrow\ket{y}$, is implemented via a quantum circuit $U_{\psi,d_{1}}({\boldsymbol{x}})$, called the quantum-embedding circuit. This circuit feeds the classical data ${\boldsymbol{x}}$ to quantum gates applied on ground states, |$\ket{0_{d_{1}}}$, in $\mathcal{H}_{d_{1}}$. To ensure computational feasibility and to avoid exponentially growing Hilbert spaces, quantum embedding circuits in the QSC framework are designed such that they compromise between the number of encoding qudits, and the dimensionality, $d_{1}$, of the Hilbert space. This optimization is mainly dependent on the size and dimensionality of the classical data, along with the available quantum computing powers. Essentially, such a quantum feature map yields the quantum embedded states |$\ket{y}$ according to: $U_{\psi,d_{1}}({\boldsymbol{x}})\ket{0_{d_{1}}}=\ket{y}$ [16]. Subsequently, these quantum states are stored in tailored structures. After embedding the data into quantum states, the next step is apply QML to construct efficient quantum semantic representations, as described next.
@@ II-B1 Problem formulation for quantum k-means clustering
where $\mathbb{H}(c\mid\ket{y})=-\sum\limits_{i}\sum\limits_{k}p(c_{k}\mid\ket{y_{i}})\ln p(c_{k}\mid\ket{y_{i}})$. The penalty term $\mathbb{I}(\ket{y};c)$ ensures that the semantic information conveyed by the $k$-means clustering that identifies the semantic concepts is maximized.
@@ II-B2 Solution to (2)
Here, the stored quantum vectors are initialized to different clusters either in an arbitrary fashion or by utilizing efficient heuristic approaches. Then, multiple iterations are performed such that in each iteration, the goal is to minimize the loss function in (2), which ensures that each vector is assigned to the cluster with the closest centroid. This process ensures that semantically similar data vectors are clustered together (as guaranteed by the penalty term in (2)). This clustering aids in uncovering hidden patterns that might remain undiscovered in the raw classical data without the utilization of . All the quantum states belonging to the same cluster form a semantic space and are semantically similar, as each centroid represents a semantic concept. Henceforth, the set $\Phi$ of |$\ket{\phi_{i}},\forall i$ forms a common language, composed of the syntactic space [13]. Along the lines of [13], the amount of semantic information conveyed by any cluster $c_{k}$ can be written as the average across of the information represented by all the elements in the cluster $c_{i}$: $\mathbb{S}(c_{k})=\sum\limits_{\ket{y}\in\mathcal{Y}}p(\ket{y})\log\frac{p(\ket{y}\mid c_{k})}{p(\ket{y})}.$ Given that the number of semantically distinct concepts $K\leq\lvert\mathcal{X}\rvert$, we chose to encode the semantic concepts into semantic representations |$\ket{s}\in\mathcal{S}$ that exists in a lower dimensional Hilbert space for quantum communication $\mathcal{H}_{d_{2}}$, where $d_{2}\leq d_{1}$. Semantic representation ensures that all the concepts that are identical from the receiver’s perspective are encoded with the same representation that gets communicated over the QCN, thus saving significant amount of quantum communication resources. A semantic representation is defined as the mapping, $f:\mathcal{Y}\,\cup\,\Phi\rightarrow\mathcal{S}$ and is implemented as a probabilistic mapping $p(\ket{s_{i}}\mid c_{k})$. The number of quantum communication resources $C=\lvert\mathcal{S}\rvert$ corresponds to the number of quantum states transmitted, which can be less than the number of clusters $K$ as the semantic representation mapping $f$ consolidates all semantically similar concepts and represents them using the same |$\ket{s_{i}}\in\mathcal{S}$.
@@ III QSC Problem Formulation and Performance Assessment Framework
First, regarding the quantum communication aspect, today’s quantum computers are devices that incorporate noise and various unavoidable losses that affect the quantum circuits employed in quantum embedding, entanglement generation, and entanglement measurement operations. A worst-case general quantum channel model that captures such noise sources in QCNs is the quantum depolarizing noise model, where a quantum state $\rho$ is preserved with a probability $1-\lambda$ and lost with probability $\lambda$, where $\lambda$ is a real-valued parameter [19]. This model, denoted as $\Delta_{\lambda}$, maps a quantum state $\rho\in\mathcal{H}_{d}$ into a linear combination that comprises $\rho$ itself, and the $d\times d$ identity matrix $\boldsymbol{I}_{d}$. It is a trace-preserving, completely positive map, given by: $\Delta_{\lambda}(\rho)=(1-\lambda)\rho+\frac{\lambda}{d}\boldsymbol{I}_{d}.$ It is important to note that parameter $\lambda$ is bounded by: $-\frac{1}{d^{2}-1}\leq\lambda\leq 1,$ so as to guarantee the complete positivity condition. The quality of the communicated quantum resources is quantified using the quantum communication fidelity, $F_{c}$, of the considered received quantum states which captures the unavoidable noise that every quantum state undergoes when transmitted over quantum channels. In this regard, when the generated entangled pairs of qudits are considered to have maximally entangled measurements and channels, the average quantum communication fidelity that corresponds for the depolarizing noise will be [19] $\langle F_{c}\rangle=1-\frac{d-1}{d}\lambda$, where $d$ is the dimension of the qudits. If a low fidelity is encountered, entanglement purification protocols may be applied [20] to achieve a desired minimum quality of the entanglement connection. Meanwhile, when the entanglement connection between the transmitter and receiver is characterized with a high fidelity, then the semantic-representing quantum states will be accurately received.